# Open-Vocabulary Segmentation with Unpaired Mask-Text Supervision

### v1

@@ 4.1 Implementation details
Datasets. During training, we randomly sample the 30% subset from the SA-1B dataset [34], which contains $\sim 3$ million images and $\sim 0.3$ billion masks. Although this supervision provides diverse binary masks, it lacks the semantic class for each mask. In addition, following Chen et al. [8], we collect about 1.3 million image-text pairs and use a large vision-language model to refine them. Afterward, we use the ChatGPT-based parser to extract entities with descriptive words from these text descriptions.
Table 1: Open-vocabulary semantic segmentation performance. We mainly compare with the fully-supervised and weakly-supervised methods. “COCO S.”, “COCO P.” and “COCO C.” denote the COCO stuff, panoptic and caption datasets. “O365” denotes the Object 365 dataset. “M. 41M” denotes the merged 41M image dataset. We report mIoU for all datasets.
@@
SimBaseline [67] COCO S. - - 15.3 - 74.5 - -
ZegFormer [16] COCO S. - - 16.4 - 73.3 - -
LSeg+ [35] COCO S. 3.8 7.8 18.0 46.5 - - -
@@
Uni-OVSeg (Ours) 30% SA1B + M. 1.3M 13.5 18.5 32.6 50.9 78.9 91.9 39.7
Training configuration. We adopt the ConvNext-Large CLIP model [43, 48] from OpenCLIP [29]. For mask segmentation training, we mainly follow Cheng et al. [13] and adopt a similar training recipe and losses without any special design. The image input size is $896\times 896$ and the batch size is 56. The model is optimised with AdamW [32, 45] with a learning rate of $1\times 10^{-4}$ and a weight decay of $0.05$. We train the model for 60k iterations and update the learning rate by a multi-step decay schedule. Besides, we adopt the same training recipe for mask-text alignment. The whole training is conducted on 56 V100-32G for $\sim 2.5$ days. $\lambda_{bce}$ and $\lambda_{dice}$ are set to 2 and 1, respectively.
Evaluation & metrics. We evaluate our model mainly on three tasks, including open-vocabulary semantic segmentation, open-vocabulary panoptic segmentation, and promptable segmentation. Following previous work [76], we adopt prompt engineering from [21, 66] and prompt templates from [22, 37]. For open-vocabulary semantic segmentation, we zero-shot evaluate the model on the COCO [40], ADE20K [84], PASCAL [18] datasets. The open-vocabulary semantic segmentation results are evaluated with the mean Intersection-over-Union (mIoU). For open-vocabulary panoptic segmentation, we evaluate the model on the COCO, ADE20K, and Cityscapes [15] datasets. We report the panoptic quality (PQ), semantic quality (SQ), and recognition quality (RQ) for open-vocabulary panoptic segmentation. For promptable segmentation, we report the 1-Point and 1-Box IoU (Oracle) on a wide range of datasets. Oracle denotes that we select the output mask with the max IoU by calculating the IoU between the prediction and target mask. More details can be found in Appendix B.
@@ 4.2 Main results
Table 2: Open-vocabulary panoptic segmentation performance. We mainly compare with the fully-supervised and unsupervised methods. “COCO P.” denotes the COCO panoptic datasets. “COCO” denotes the COCO image dataset. “IN 1K” denotes the ImageNet-1K image dataset. We report PQ, SQ and RQ for all datasets.
@@
Uni-OVSeg (Ours) 30% SA1B + M. 1.3M 18.0 72.6 24.3 14.1 66.1 19.0 17.5 65.2 23.5
@@
Figure 3: Point-promptable segmentation performance. We compare our method with SAM-ViT/L [34] on a wide range of datasets. Given a $20\times 20$ point grid as visual prompt, we select the output masks with max IoU by calculating the IoU with the ground-truth masks. We report 1-pt IoU for all datasets.
Open-vocabulary semantic segmentation. As demonstrated in Tab. 1, our study presents a comprehensive comparison of our Uni-OVSeg against previous works across a range of benchmarks. These include ADE20K (encompassing both 150 and 847 class variants), PASCAL Context (459 and 59 class variants), PASCAL VOC (with 20 and 21 class categories), and Cityscapes. Compared to weakly-supervised methods, Uni-OVSeg exhibits remarkable performance improvements across all evaluated datasets. Specifically, in the more challenging datasets of PASCAL Context-459, Uni-OVSeg not only surpasses its weakly-supervised counterparts but also outperforms the cutting-edge fully-supervised methods, e.g., FC-CLIP. This is indicative of Uni-OVSeg’s superior capability in categorizing a diverse array of semantic classes. Furthermore, in the PASCAL VOC benchmarks (20 and 21 classes), Uni-OVSeg demonstrates a substantial enhancement over state-of-the-art weakly-supervised methods, achieving improvements of 18.3% and 12.2% mIoU, respectively, which demonstrates our Uni-OVSeg captures fine-grained spatial structures. These results elevate the practical applicability of weakly-supervised open-vocabulary segmentation to new heights.
Open-vocabulary panoptic segmentation. For open-vocabulary panoptic segmentation, we zero-shot evaluate our model on the COCO, ADE20K, and Cityscapes datasets, shown in Tab. 2. Existing weakly-supervised methods only use text supervision, which causes a challenge to discriminate different instances with the same semantic class. To the best of our knowledge, we are the first to learn open-vocabulary panoptic segmentation with weak supervision. Compared to unsupervised methods, we obviously outperform U2Seg by 1.9% PQ, 1.5%SQ, and 4.4% RQ on COCO datasets. Unfortunately, our used image-mask pairs contain multiple granularity masks, such as object-wise and part-wise masks, which is different with the panoptic segmentation datasets. This discrepancy leads to a number of false positive results, limiting our performance.
Promptable segmentation. To evaluate the segmentation efficacy of Uni-OVSeg in scenarios involving interactive point and box prompts, we conduct comparative analyses with the SAM-ViT/L model across a range of datasets from diverse domains. For visual prompts, we implement a uniform 20$\times$20 point grid as the interactive point prompt and utilise the actual bounding boxes as box prompts. The segmentation performance is measured using the 1-pt IoU (Oracle) metric across all datasets. The results, as demonstrated in Fig. 3, indicate that Uni-OVSeg exceeds SAM in most datasets when employing point and box prompts. In particular, Uni-OVSeg significantly surpasses SAM in terms of IoU by 31.1%, 20.1%, and 43.2% on the TimberSeg, COCO instance, and PIDray datasets, respectively. Additionally, we provide a visualisation of the segmentation results of both SAM and Uni-OVSeg in Fig. 4. The visualisation reveals that SAM tends to miss larger masks and erroneously categorise areas with unclear boundaries. In contrast, Uni-OVSeg demonstrates exceptional precision in segmenting objects of varying scales and different stuff. More experimental results are visualisation can be found in the Appendix B and C.
Figure 4: Visualisaton of point-promptable automatic mask generation. We adopt a $20\times 20$ point grid as a visual prompt and select the output masks with max IoU by calculating the IoU with the ground truth masks.
Table 3: Ablation study on mask-text alignment.Refine.” denotes the text refinement by the LVLM. “Parser.” denotes the text parser, which extracts entities from text descriptions. “NLTK” and “GPT” denote the natural language toolkit and ChatGPT-based parser. “M.S.” denotes the multi-scale ensemble strategy.
Table 4: Mask classification performance. we first pool the region features based on the provided ground truth masks. These pooled features are then projected into the CLIP embedding space, where they are classified using text embeddings. We report the Top-1 accuracy (%) and time (sec. / sample).
Refine. Parser. M.S. COCO PC-59
PQ (%) mIoU (%) mIoU (%)
baseline 13.2 33.2 39.7
NLTK 14.1 34.5 40.9
NLTK 15.6 37.3 44.5
GPT 16.2 40.4 48.2
GPT 18.0 42.7 50.9
Input size ADE20K COCO Time
768 45.2 49.0 0.10
896 48.3 53.4 0.13
1024 50.0 55.5 0.16
768, 896 49.2 55.2 0.20
896, 1024 51.3 58.3 0.26
768, 896, 1024 51.6 58.4 0.35
Table 4: Mask classification performance. we first pool the region features based on the provided ground truth masks. These pooled features are then projected into the CLIP embedding space, where they are classified using text embeddings. We report the Top-1 accuracy (%) and time (sec. / sample).
@@ 4.3 Ablation study
We conduct an extensive ablation study to demonstrate the contribution of each component of our framework.
Mask-text alignment. Compared to the straightforward baseline, as shown in Tab. 4, our proposed Uni-OVSeg achieves significant gains of 4.8% PQ and 9.5% mIoU on the COCO dataset, and 11.2% mIoU on the PASCAL Context-59 dataset. This demonstrates our method effectively align objects in images and entities in text descriptions, generalising the CLIP embedding space from the image level to pixel level. By resorting to the refinement of text descriptions, new texts are more correlated with the corresponding images, improving the mIoU from 34.5% to 37.3% on the COCO dataset. Compared to the traditional NLP toolkit (NLTK) [3], ChatGPT-based parser extracts more reliable entities from text descriptions, which achieves obvious improvements of 3.1% and 3.7% mIoU on the COCO and PASCAL Context-59 datasets, respectively. Finally, the proposed multi-scale ensemble strategy that leverages the multi-scale information of objects within the images, stabilise the mask-text matching, which achieves a performance gain of 1.8% PQ on the COCO datasets.
Multi-scale ensemble in mask-text matching. The quality of correspondence between masks and entities is an essential part of mask-text matching. To investigate the impact of multi-scale information on this correspondence, as illustrated in Tab. 4, we use masks and semantic classes from the ADE20K and COCO datasets, reporting the Top1 accuracy and forward time per sample. We first resize input images to multiple resolutions and extract visual features via the clip visual encoder. Given ground-truth masks, regional features are pooled from CLIP visual features and projected into the clip embedding space. Each regional embedding is classified by text embeddings. Taking into account the trade-off between performance and latency, we adopt the sizes of $869\times 896$ and $1024\times 1024$ as default.
@@ 5 Conclusion
In conclusion, this paper proposes an innovative framework for weakly-supervised open-vocabulary segmentation, named Uni-OVSeg. Using independent image-text and image-mask pairs, Uni-OVSeg effectively reduces the dependency on labour-intensive image-mask-text triplets, meanwhile achieving impressive segmentation performance in open-vocabulary settings. Using the LVLM to refine text descriptions and multi-scale ensemble to enhance the quality of region embeddings, we alleviate the noise in mask-text correspondences, achieving substantial performance improvements. Notably, Uni-OVSeg significantly outperforms previous state-of-the-art weakly-supervised methods and even surpasses the cutting-edge fully-supervised method on the Challenging PASCAL Context-459 dataset. This impressive advancement demonstrates the superiority of our proposed framework and paves the way for further research.
6 Broader impacts
The Uni-OVSeg framework represents a significant advancement in open-vocabulary segmentation by reducing the dependency on labour-intensive image-mask-text triplet annotations. This innovation has the potential to democratise access to cutting-edge vision perception systems, offering substantial benefits across various sectors, such as medical imaging and autonomous vehicles. The development of a more efficient and accurate vision perception system contributes to the community, potentially leading to more innovative applications and research in machine learning, computer vision, and related areas. As with any AI model, there is a risk of bias in the data used for training. Efforts must be made to ensure that the datasets are diverse and representative to avoid perpetuating or amplifying biases.
@@ Appendix A Framework details
Visual prompt encoder. We have two types of visual prompts, including points and bounding boxes. Each type of prompts is mapped to 256-wide embeddings as follows. A point is first converted into a small bounding box. We represent a bounding box by an embedding pair, including the sine position embedding of its top-left and bottom-right corners. Afterward, we use two extra learnable embeddings to indicate these two corners.
@@
Mask decoder. Following Mask2Former, we use a similar transformer decoder design. Each visual prompt embedding within our framework is coupled with a sine positional embedding. As depicted in Fig. 5, our approach utilizes six mask decoder layers, and we apply the same loss function after each layer. Moreover, every decoder layer includes a cross-attention layer. This layer is crucial as it ensures that the final pixel features are enriched with essential geometric information, such as point coordinates and bounding boxes. Besides, in each decoder layer, we use a cross-attention layer, which ensures that the final pixel features have access to critical geometric information (e.g., point coordinates, and boxes). Finally, visual prompt embeddings are fed to a layer normalization, followed by processing through a multiple perception layer (MLP). These processed embeddings are then dot-producted with the pixel features, culminating in the generation of mask predictions.
@@
Figure 6: Architecture of the multi-scale feature adapter. Each LoRA adapter is used to update a specific downscaled features.
@@
Multi-scale feature adapter. The multi-scale feature adapter illustrated in Fig. 6 is designed to refine and enhance pixel features at various scales. This adapter is composed of multiple Low-rank Adapters (LoRA adapters) [27], each specifically tailored to update features at a particular scale. Each LoRA adapter incorporates two linear layers, denoted as A and B, with a non-linear activation function placed in between. The linear layers A and B serve to transform the input data linearly, while the activation function introduces non-linearity, allowing for the modeling of more complex relationships in the data. Each adapter is responsible for handling features at a specific scale, suggesting a hierarchical approach to feature refinement. This modular design allows for a focused and specialized treatment of features depending on their resolution and semantic complexity, which can be particularly advantageous for dense prediction tasks.
Figure 7: Box-promptable segmentation performance. We compare our method with SAM-ViT/L [34] on a wide range of datasets. Given a ground-truth box as the visual prompt, we select the output masks with max IoU by calculating the IoU with the ground-truth masks. We report 1-pt IoU for all datasets.
@@
Figure 8: Point-promptable segmentation performance. We compare our method with SAM-ViT/L [34] on the SegInW datasets [87]. Given a $20\times 20$ point grid as a visual prompt, we select the output masks with max IoU by calculating the IoU with the ground-truth masks. We report 1-pt IoU for all datasets.
@@
Figure 9: Box-promptable segmentation performance. We compare our method with SAM-ViT/L [34] on the SegInW datasets [87]. Given a ground-truth box as the visual prompt, we select the output masks with max IoU by calculating the IoU with the ground-truth masks. We report 1-pt IoU for all datasets.

### v2

@@ 4.1 Implementation details
Datasets. During training, we randomly sample the 50% subset from the SA-1B dataset [32], which contains $\sim 5$ million images and $\sim 0.5$ billion masks. Although this supervision provides diverse binary masks, it lacks the semantic class for each mask. Besides, we collect about 130k images [3, 13] and adopt a vision-language large model [16] to generate captions, forming image-text pairs. Afterwards, we extract entities from these text descriptions.
Table 1: Open-vocabulary semantic segmentation performance. We mainly compare with the fully-supervised and weakly-supervised methods. “COCO S.”, “COCO P.”, and “COCO C.” denote the COCO stuff, panoptic, and caption datasets. “O365” denotes the Object 365 dataset. “M. 41M” denotes the merged 41M image dataset. We report mIoU for all datasets.
@@
LSeg [33] COCO S. 3.8 7.8 18.0 46.5 - - -
@@
Baseline (Ours) 50% SA1B 12.1 13.3 27.0 40.6 60.3 90.6 27.2
Unpair-Seg (Ours) 50% SA1B + M. 130K 14.6 19.5 32.6 52.2 76.1 93.0 45.2
Training configuration. We adopt our visual and text encoders as our ConvNext-Large CLIP model [41, 46] from OpenCLIP [27]. For mask segmentation training, we mainly follow [11] and adopt a similar training recipe and losses without any particular design. The input size of the image is $1024\times 1024$, and we set the batch size to 64. The model is optimised with AdamW [30, 43], where the learning rate is set to $1\times 10^{-4}$ and the weight decay is set to $5\times 10^{-2}$. We train the model for 200k iterations and update the learning rate via the multi-step decay schedule. The parameters $\lambda_{bce}$, $\lambda_{dice}$, and $\lambda_{iou}$ are set to 2, 1, and 5, respectively. Afterwards, during the training process of mask-entity alignment, we adopt a $20\times 20$ point grid as the input visual prompt. The batch size and learning rate are set to 32 and $2\times 10^{-5}$, respectively. We train this stage for 30k iterations. We conduct all training on 8 NVIDIA H100 80GB, 109 hours for the first stage, and 7 hours for the second stage.
Evaluation & metrics. We evaluate our model mainly on four tasks: open-vocabulary semantic segmentation, open-vocabulary panoptic segmentation, point-prompt segmentation, and box-prompt segmentation. Following previous work [71], we adopt prompt engineering from [20, 62] and prompt templates from [21, 35]. For open-vocabulary semantic segmentation, we zero-shot evaluate the model on the COCO [38], ADE20K [79], and PASCAL [17] datasets. Following previous work [71], we adopt the prompt engineering from [20, 62, 21, 35]. The open-vocabulary semantic segmentation results are evaluated with the mean Intersection-over-Union (mIoU). We evaluate the model on the COCO, ADE20K, and Cityscapes [13] datasets for open-vocabulary panoptic segmentation. We report the panoptic quality (PQ), semantic quality (SQ), and recognition quality (RQ) for open-vocabulary panoptic segmentation. For point-prompt and box-prompt segmentation, from the oracle perspective, we report the 1-pt IoU and 1-box IoU on a wide range of datasets. “Oracle” denotes that we select the output mask with the max IoU by calculating the IoU between the prediction and ground-truth mask. More details can be found in Appendix A and Appendix C.
@@ 4.2 Main results
Open-vocabulary semantic segmentation. To showcase the significant reduction in the performance gap between fully-supervised and weakly-supervised methods, we conducted a comprehensive comparison across various datasets, including ADE20K (encompassing both 150 and 847 class variants) [79], PASCAL Context (459 and 59 class variants) [17], PASCAL VOC (with 20 and 21 class categories) [17], and Cityscapes [13]. As shown in Table 1, Unpair-Seg outperforms weakly-supervised methods across different datasets by utilising independent image-mask and image-text pairs. Notably, Unpair-Seg performs impressive on the challenging PASCAL Context-459 and ADE20K-847 datasets, surpassing the state-of-the-art fully-supervised method FC-CLIP [71]. Upon further analysis, we observed that FC-CLIP achieves significantly higher accuracy for in-vocabulary classes (corresponding to the COCO dataset) than out-of-vocabulary ones on the PASCAL Context-459 and ADE20K-847 datasets. More details can be found in Appendix B. Leveraging the diversity of image-text pairs, our method demonstrates superior capability in categorising general semantic classes.
Table 2: Open-vocabulary panoptic segmentation performance. We mainly compare with the fully-supervised and unsupervised methods. “COCO P.”, “COCO” and “IN 1K” denote the COCO panoptic, COCO image and ImageNet-1K datasets, respectively. We report PQ, SQ and RQ for all datasets. Fully-supervised method trained on the “COCO P.” datasets, so we show the results in grey.
@@
Baseline (Ours) 50% SA1B 24.4 79.6 30.0 19.1 78.1 23.7 17.3 71.4 22.0
Unpair-Seg (Ours) 50% SA1B + M. 130K 28.8 80.0 35.2 22.6 74.0 28.0 26.8 70.9 33.4
@@
Table 3: Promptable segmentation performance. We compare our method with SAM [32]. Given point or box prompts, we select the output masks with max IoU by calculating the IoU with the ground-truth masks. We report 1-pt IoU on the COCO and ADE20K panoptic segmentation, and 1-box IoU on the COCO and ADE20K instance segmentation datasets, respectively.
Method Param Training Data COCO ADE20K
1-pt IoU (%) 1-box IoU (%) 1-pt IoU (%) 1-box IoU (%)
SAM-Base 93M 100% SA-1B 42.7 72.8 41.6 74.7
SAM-Large 312M 100% SA-1B 64.3 77.9 61.1 77.7
SAM-Huge 641M 100% SA-1B 66.7 78.0 63.3 78.3
Unpair-Seg (Ours) 221M 50% SA-1B 80.8 76.9 76.9 77.3
Open-vocabulary panoptic segmentation. Existing weakly-supervised methods only utilise text supervision, which makes it challenging to distinguish different instances with the same semantic class. To the best of our knowledge, Unpair-Seg is the first attempt to tackle the difficult task of open-vocabulary panoptic segmentation with weak supervision. Due to the absence of benchmarks, we evaluate our method on the COCO, ADE20K, and Cityscapes datasets and compared it comprehensively with fully-supervised and unsupervised methods in Table 2. Compared to unsupervised methods, our approach outperforms U2Seg [45] by 12.7% PQ, 8.9% SQ, and 15.3% RQ on the COCO panoptic segmentation task. Even more impressively, Unpair-Seg exceeds specific fully-supervised methods. For instance, we achieve a noticeable gain of 3.6% PQ than OPSNet [8] on the ADE20K dataset and a remarkable improvement of 2.9% PQ than ODISE [62] on the Cityscapes panoptic segmentation. These results show that our method effectively captures fine-grained spatial relations through weak supervision, thereby enhancing the practical usefulness of weakly-supervised open-vocabulary segmentation.
Promptable segmentation. Following SAM [32], we evaluate segmenting an object from one input point, which is challenging as one point can refer to objects at multiple levels of detail. To provide visual cues, we have implemented a uniform $20\times 20$ point grid as the interactive point prompt and use the actual bounding boxes as box prompts. We compare our method with three variants of SAM, as shown in Table 3. By using the proposed many-to-many matching, our method can predict multi-granularity masks when employing point prompts. In particular, Unpair-Seg significantly outperforms SAM-Large in terms of 1-point IoU by 16.5% and 15.8% on the COCO and ADE20K panoptic segmentation datasets, respectively. We also show some visualisations in Figure 4 and more experimental results and visualisations can be found in Appendix C and E.
Table 4: Vision-language large model.LLaVA.” and “InternLM-VL.” denote the LLaVA-v1.5-7B and InternLM-XComposer2-VL-7B models.
Method COCO PC-59
PQ (%) mIoU (%)
LLaVA. 25.4 42.6
InternLM-VL. 28.8 52.2
Table 5: Design choice of the feature adapter.Cross Attn.” and “Low Rank.” denote the cross attention-based and low rank-base feature adapters.
Method COCO PC-59
PQ (%) mIoU (%)
Cross Attn. 27.1 47.4
Low Rank. 28.8 52.2
Table 6: Mask-entity matching strategy.Greedy.” and “Bipartite.” denote the confidence greedy and bipartite match strategies.
Method COCO PC-59
PQ (%) mIoU (%)
Greedy. (1 sigma) 26.1 43.7
Greedy. (2 sigma) 26.0 43.6
Bipartite. 28.8 52.2
Table 7: Multi-scale ensemble. We ensemble multi-scale visual embeddings extracted by the CLIP model to stablise the matching process.
Size COCO PC-59
PQ (%) mIoU (%)
896 27.4 49.0
896, 1024 28.2 50.0
896, 1024, 1152 28.8 52.2
Figure 4: Visualisaton. We show prediction results on three tasks: promptable, open-vocabulary semantic, and open-vocabulary panoptic segmentation. The results are best viewed in color.
@@ 4.3 Ablation study
We conducted a thorough ablation study to assess the impact of the main components in our framework. Compared to the simple baseline, Unpair-Seg achieved significant improvements of 2.5%, 6.2%, and 18.0% mIoU on the ADE20k-847, PASCAL Context-459, and Cityscapes semantic segmentation, and 4.4% and 3.5% PQ on the COCO and ADE20K panoptic segmentation, as shown in Table 1 and Table 2. These results demonstrate the effectiveness of our method in aligning objects in images and entities in text descriptions, generalising the CLIP embedding space from the image to the pixel level. Qualitatively, we illustrate a visualisation comparison between the baseline and our proposed Unpair-Seg in Figure 4. Furthermore, we explored the impact of different vision-language large models on the quality of image-text pairs. Our ablation study on two models, LLaVA [40] and InternLM-VL [16], detailed in Table 5, revealed their influence on model performance. High-quality captions were found to help the model learn accurate mask-entity alignment. During the mask-entity alignment, we use a feature adapter to enhance regional embeddings, followed by aligning with text embedding. In Table 5, we design two alternative modules to aggregate multi-scale visual information and improve the quality of regional embedding. A low-rank feature adapter that is adopted in our framework achieves better results. As illustrated in Table 7, we design two mask-entity matching strategies. The greedy method assigns an entity for each predicted mask based on the maximum confidence. The Bipartite method optimises an assignment plan to achieve the global minimum cost, which obtains better results than the greedy method. Finally, considering the variety of object scales, we employ a multi-scale ensemble that extracts CLIP visual embeddings from multiple resolution images and ensemble them into one embedding, which is used to obtain confident mask-entity pairs. As shown in Table 7, we adopt the sizes of $869\times 896$, $1024\times 1024$, and $1152\times 1152$ as default.
@@ 5 Conclusion
This paper introduces Unpair-Seg, a novel framework for weakly-supervised open-vocabulary segmentation. The key strength of Unpair-Seg lies in its use of unpaired mask-text supervision, which reduces the need for labor-intensive annotations. To mitigate inherent noise in the mask-entity matching process, we employ a vision-language large model for precise entity extraction and a multi-scale matching strategy for stable mask-entity alignment. By predicting binary masks and generating pseudo labels through confident mask-text entity pairs, Unpair-Seg significantly outperforms previous weakly-supervised methods across multiple segmentation tasks and datasets. This significant improvement brings us closer to the performance of fully-supervised methods in real-world scenarios. However, it’s important to acknowledge that the framework’s performance may be influenced by the quality of image-text pairs and initial mask predictions in complex or ambiguous scenes. These aspects will be the focus of our future work.
@@ Appendix A Framework details
Visual prompt encoder. We have two types of visual prompts, including points and bounding boxes. Each type of prompt is mapped to 256-wide embeddings as follows. A point is first converted into a small bounding box. We represent a bounding box by an embedding pair, including the sine position embedding of its top-left and bottom-right corners. Afterward, we use two extra learnable embeddings to indicate these two corners.
@@
Mask decoder. Following Mask2Former, we use a similar transformer decoder design. Each visual prompt embedding within our framework is coupled with a sine positional embedding. As depicted in Figure 5, our approach utilizes six mask decoder layers, and we apply the same loss function after each layer. Moreover, every decoder layer includes a cross-attention layer. This layer is crucial as it ensures that the final pixel features are enriched with essential geometric information, such as point coordinates and bounding boxes. Besides, in each decoder layer, we use a cross-attention layer, which ensures that the final pixel features have access to critical geometric information (e.g., point coordinates, and boxes). Finally, visual prompt embeddings are fed to a layer normalization, followed by processing through a multiple perception layer (MLP). These processed embeddings are then dot-producted with the pixel features, culminating in the generation of mask predictions.
@@
Figure 6: Architecture of the multi-scale feature adapter. Each low-rank feature adapter is used to update specific downscaled features.
@@
Multi-scale feature adapter. The multi-scale feature adapter illustrated in Figure 6 is designed to refine and enhance pixel features at various scales. This adapter is composed of multiple Low-rank Adapters [26], each specifically tailored to update features at a particular scale. Each feature adapter incorporates two linear layers, denoted as A and B, with a non-linear activation function placed in between. The linear layers A and B serve to transform the input data linearly, while the activation function introduces non-linearity, allowing for the modeling of more complex relationships in the data. Each adapter is responsible for handling features at a specific scale, suggesting a hierarchical approach to feature refinement. This modular design allows for focused and specialized treatment of features depending on their resolution and semantic complexity, which can be particularly advantageous for dense prediction tasks.
@@
Figure 7: Point-prompt segmentation performance. We compare our method with SAM-ViT/L [32]. Given a $20\times 20$ point grid as visual prompt, we select the output masks with max IoU by calculating the IoU with the ground-truth masks. We report 1-pt IoU for all datasets.
@@
Figure 8: Box-prompt segmentation performance. We compare our method with SAM-ViT/L [32]. Given a ground-truth box as the visual prompt, we select the output masks with max IoU by calculating the IoU with the ground-truth masks. We report 1-pt IoU for all datasets.
Figure 9: Point-prompt segmentation performance on the SegInW dataset. We compare our method with SAM-ViT/L [32]. Given a $20\times 20$ point grid as a visual prompt, we select the output masks with max IoU by calculating the IoU with the ground-truth masks. We report 1-pt IoU for all datasets.
Figure 10: Box-prompt segmentation performance on the SegInW dataset. We compare our method with SAM-ViT/L [32]. Given a ground-truth box as the visual prompt, we select the output masks with max IoU by calculating the IoU with the ground-truth masks. We report 1-box IoU for all datasets.
Table 8: In- and out-vocabulary mIoU on ADE20K-847 semantic segmentation. We report the results of FC-CLIP and our method, respectively.
Method In. (101 classes) Out. (746 classes)
FC-CLIP [71] 35.1 12.0
Unpair-Seg 32.5 12.1
Table 9: In- and out-vocabulary mIoU on PASCAL Context-459 semantic segmentation. We report the results of FC-CLIP and our method, respectively.
Method In. (85 classes) Out. (374 classes)
FC-CLIP [71] 47.5 9.5
Unpair-Seg 42.4 12.7
Appendix B In- & Out-vocabulary IoU comparison
As shown in Table 9 and Table 9, we report in-vocabulary and out-vocabulary IoU on ADE20K-847 and PASCAL Context-459 datasets. FC-CLIP, trained on the COCO panoptic dataset in a fully supervised manner, exhibits higher mIoU for in-vocabulary classes (35.1% on ADE20K-847 and 47.5% on PASCAL Context-847), suggesting a tendency to overfit on these classes, thus limiting its open-vocabulary segmentation ability. In contrast, Unpair-Seg, which leverages the text diversity from large amounts of image-text pairs, demonstrates competitive in-vocabulary performance (32.5% on ADE20K-847 and 42.4% on PASCAL Context-847, respectively) and significantly better out-of-vocabulary mIoU (12.1% and 12.7%), indicating superior generalisation to unseen classes and robustness in sophisticated scenarios.
Appendix C Promptable segmentation
Evaluation details. We perform a prompt segmentation evaluation on a wide range of datasets in various domains. For the point prompt, we adopt a uniform point grid $h\times w$ as input prompts (e.g., 20 $\times$ 20). For the box prompt, we use ground-truth bounding boxes as input prompts. 1-pt IoU denotes the oracle performance of one point by evaluating the intersection-over-union (IoU) of the predicted masks that best match ground truth. 1-box IoU denotes is similar to 1-pt IoU. More evaluation results are reported in Figure 7, Figure 8, Figure 9 and Figure 10.